# Action-Driven Flows for Causal Variational Principles

### v1

@@ 1. Introduction
In the physical theory of causal fermion systems, spacetime and the structures therein are described by a minimizer of the so-called causal action principle (for an introduction to the physical background and the mathematical context, we refer the interested reader to the reviews [9, 11], the textbooks [7, 10] or the website [1]). Causal variational principles evolved as a mathematical generalization of the causal action principle [6, 12]. The goal of this paper is to develop action-driven flows for causal variational principles and the causal action principle. Our flow is a non-smooth version of the gradient flow inspired by De Giorgi’s minimizing movements approach [5, 2, 4]. In order to develop our methods and ideas step by step, we begin with a few general remarks on non-smooth and non-convex variational problems (Section 3). The core of the paper is to develop the method of minimizing movements for causal variational principles in the simplest so-called compact setting, where the underlying space $\mathscr{F}$ is assumed to be compact (Section 4). Then we adapt and generalize these methods to the causal action principle in finite dimensions (Section 6). Here the main technical tool is to work with moment measures as introduced in [6, Section 2] (see also [10, Chapter 12]). As an application, we work with a filtration by finite-dimensional subspaces to construct a flow for the causal action principle in infinite dimensions (Section 7).
We now explain our methods and results in more detail in the simplest case of causal variational in the compact setting (for more details see the preliminaries in Section 2.1). Therefore, our starting point is a compact metric space $\mathscr{F}$ and a non-negative function :×${{\mathcal{L}}}:\mathscr{F}\times\mathscr{F}\rightarrow\mathbb{R}^{+}_{0}:=[0,\infty)$ (the Lagrangian) which is assumed to be continuous. The corresponding causal action principle then is to
@@
over the class $\mathfrak{M}_{1}(\mathscr{F})$ of normalized Borel measures on $\mathscr{F}$. Causal variational principles are a class of examples for non-smooth and non-convex variational principles.
Minimizers can be constructed with the direct method of the calculus of variations (see the preliminaries in Section 2.1). A minimizer $\rho$ satisfies corresponding Euler-Lagrange (EL) equations (see (2.1) in Section 2.1). Constructing solutions of the EL equations is of central importance in the theory of causal fermion systems for understanding the nature of the physical interactions as described by the causal action principle. Here abstract existence results are not sufficient, but one needs constructive methods which give insight into the structure of the minimizing measure. This is our main motivation for the present paper, where we shall construct a flow
@@
with the property that the action is strictly decreasing along the flow lines. We begin with the flow obtained by minimizing movements, where as the distance function used for the penalization we choose either the Wasserstein metric or the total variation norm. In this way, we obtain a Hölder continuous curve of measures $\varrho(t)$. If this curve converges, we obtain a desired solution of the Euler-Lagrange equations. We remark that the smooth and finite-dimensional analog of our flow is the ordinary gradient flow described by the ordinary differential equation
$\frac{\operatorname{d}\!}{\operatorname{d}\!t}\varrho(t)=-\nabla{\mathcal{S}}(\varrho)\>,$
with the gradient defined via the metric on $\mathscr{F}$. In our general setting, however, this equation makes no mathematical sense, simply because the derivative on the left in general does not exist (apart from the fact that, since the space of signed measures is infinite-dimensional, one would have to introduce the derivatives properly for example as Fréchet derivatives). Minimizing movements can be understood as a method for extending the gradient flow to non-smooth actions on infinite-dimensional spaces.
@@
A drawback of this flow is that the curve $\varrho(t)$ will in general not converge in the limit $t\rightarrow\infty$ (as one already sees in the two-dimensional example of a non-smooth non-convex variational principle discussed in Section 3). Moreover, even if the curve does converge, the above flow may not be helpful for practical applications or numerical studies because we have no control on how long it takes to reach an approximate solution of the EL equations. In order to improve the situation, we also introduce another flow which involves an additional penalization term involving a parameter $\xi>0$. In the case $\xi=0$, we get back the above flow by minimizing movements. In the case $\xi>0$, however, we prove that the resulting curve $\varrho^{\xi}(t)$ does converge and that the length of this curve is bounded a-priori by ${\mathcal{S}}(\varrho^{\xi}(0))/\xi$. The prize to pay is that the limiting measure satisfies the EL equations only approximately. For the error term, we derive a precise a-priori bound which tends to zero as $\xi\searrow 0$. With this in mind, our procedure seems well-suited for the applications in mind. For example, in a numerical study one can choose $\xi$ so small that the error of the approximation is bounded by the numerical errors.
Our methods and results can be understood more generally from the perspective of non-convex variational problems. Indeed, causal variational principles are model examples of variational principles which, in general, are fully non-convex. Different from other settings which deal with convex scenarios (see, e.g.,  [3]), the non-convexity implies that the gradient flow constructed by the minimizing movement scheme does not necessarily converge to a minimizer of (1.1). In fact, by the very definition of the flow, it might get stuck at a critical point of the functional, and by the non-convexity, the latter might be far away from any global minimizer. It is thus reasonable to strive for a convergence result for the flow as $t\to\infty$ which merely makes assertions on the validity of the Euler-Lagrange equations and not minimality. However, in the general case considered here, it is again the non-convexity of (1.1) which rules out this possibility: As we explain in Section 3 by use of a particular example of an action $\mathcal{S}$ (constructed as a downward spiral with increasingly small potential wells), which might be known to the experts in different scenarios, it might happen that the gradient flow does not converge at all. Even more, there is not even a subsequence of times $(t_{k})$ for which the measures converge to a solution of the EL equations. Therefore, in order to quantitatively capture the behavior of the flows, it is reasonable to introduce approximations of the original variational problem which enforce convergence of the associated flows and give control of the length of the curves. The methods to be developed in the present paper provide Hölder continuous flows of measures with these desired properties.
@@ 2.1. Causal variational principles in the compact setting
We let $\mathscr{F}$ be a compact topological space and suppose that the Lagrangian :×${\mathcal{L}}\colon\mathscr{F}\times\mathscr{F}\to\mathbb{R}_{0}^{+}$ satisfies the following assumptions:
@@
    ${\mathcal{L}}\in\operatorname{C}^{0}(\mathscr{F}\times\mathscr{F},\mathbb{R}^{+}_{0})$ is continuous in both arguments.
@@
${\mathcal{S}}(\rho)=\int_{\mathscr{F}}\operatorname{d}\!\rho(x)\int_{\mathscr{F}}\operatorname{d}\!\rho(y)\>{\mathcal{L}}(x,y)$
@@
$M:=\operatorname{supp}\rho:=\mathscr{F}\setminus\bigcup\big{\{}U\subset\mathscr{F}\;\text{open}\colon\;\rho(U)>0\big{\}}\>.$
@@ 2.2. Background facts from optimal transport and metric measure spaces
In our arguments below, we will also make use of the $p$-Wasserstein metric on $X:=\mathfrak{M}_{1}(\mathscr{F})$ for $1\leqslant p<\infty$. Given a measure $\mathbb{P}\in\mathfrak{M}_{1}(X\times X)$, for $i\in\{1,2\}$ we denote the projection to the $i^{\text{th}}$ component by $\pi^{i}\colon X\times X\ni(x_{1},x_{2})\mapsto x_{i}\in X$. We let $\pi_{\#}^{i}\mathbb{P}(A):=\mathbb{P}(\pi_{i}^{-1}(A))$ be the corresponding push-forward of $\mathbb{P}$. As is customary in this context, we then define for $\mu_{1},\mu_{2}\in\mathfrak{M}_{1}(\mathscr{F})$ the class of couplings $\Gamma(\mu_{1},\mu_{2})$ (also referred to as transport plans) by
@@
Let $1\leqslant p<\infty$. We then define for $\mu,\nu\in\mathfrak{M}_{1}(\mathscr{F})$ the $p$-th Wasserstein metric by
@@
It has the desired marginals. We thus obtain the estimate
[…]
@@
$=\text{diam}(\mathscr{F})^{p}\>\|\mu-\nu\|_{\mathfrak{M}(\mathscr{F})}\>.$
@@
In order to prove (2.6), we let $\varepsilon>0$ be arbitrary and choose $\mathbb{P}\in\Gamma(\mu,\nu)$, ~$\widetilde{\mathbb{P}}\in\Gamma(\nu,\nu)$ such that
@@ 3. An example of a non-smooth, non-convex variational principle
$\dot{\gamma}(t)=-\nabla{\mathcal{S}}|_{\gamma(t)}\qquad\text{and}\qquad\gamma(0)=\Big{(}r=\frac{1}{5},\varphi=0\Big{)}\>.$
@@
${\mathcal{S}}(r,\varphi)+\frac{1}{2h}\>d\big{(}(r,\phi),(r^{\prime},\phi^{\prime})\big{)}^{2}\>,$
where $d$ denotes the Euclidean distance in $\mathbb{R}^{2}$.
@@
The corresponding flow equation takes the form
@@ 4.1. The causal action with penalization
[…]
@@
Let $\mathscr{F}$ be a compact topological space and let ${\mathcal{L}}\in\operatorname{C}(\mathscr{F}\times\mathscr{F})$. Then the functional
@@
If we assume that the Lagrangian ${\mathcal{L}}\in\operatorname{C}^{0,\alpha}(\mathscr{F}\times\mathscr{F},\mathbb{R}^{+}_{0})$ is Hölder continuous with Hölder exponent $\alpha\in(0,1]$, then so is the functional ${\mathcal{S}}$ with respect to the Wasserstein distance, i.e. there is a constant $C$ (which again depends only on $\mathscr{F}$ and ${\mathcal{L}}$) such that for all $\rho,\tilde{\rho}\in\mathfrak{M}_{1}(\mathscr{F})$,
@@
Let $\rho,\rho_{1},\rho_{2},...\in\mathfrak{M}(\mathscr{F})$ be such that $\rho_{j}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\rho$ as $j\to\infty$. By the Banach-Steinhaus theorem, $(\rho_{j})$ is bounded in $\mathfrak{M}(\mathscr{F})$: $m:=\sup_{j\in\mathbb{N}}\|\rho_{j}\|_{\mathfrak{M}(\mathscr{F})}<\infty$. Since $\mathcal{F}$ is compact, the Weierstraß  approximation theorem implies that the space
@@
is dense in $\operatorname{C}(\mathscr{F})$. Let $\varepsilon>0$ be arbitrary but fixed. We then find $h\in X$, $h(x,y)=\sum_{i=1}^{N}h_{i}f_{i}(x)g_{i}(y)$, $h_{1},...,h_{N}\in\mathbb{R}$, such that $\|{\mathcal{L}}-h\|_{\infty}<\varepsilon$. Therefore,
|×$\left|\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)-\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)\right|$
×$\leqslant\iint_{\mathscr{F}\times\mathscr{F}}|{\mathcal{L}}(x,y)-h(x,y)|\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)$
+|×$\quad\>+\left|\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\,(\rho_{j}-\rho)(x)\operatorname{d}\!\,(\rho_{j}-\rho)(y)\right|$
+×$\quad\>+\iint_{\mathscr{F}\times\mathscr{F}}|{\mathcal{L}}(x,y)-h(x,y)|\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)=:\mathrm{I}+\mathrm{II}+\mathrm{III}.$
@@
×$\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)$ $=\sum_{i=1}^{N}h_{i}\Big{(}\int_{\mathscr{F}}f(x)\operatorname{d}\!\rho_{j}(x)\Big{)}\Big{(}\int_{\mathscr{F}}g(y)\operatorname{d}\!\rho_{j}(y)\Big{)}$
@@
=×$=\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)$
@@
Since :×${\mathcal{L}}\colon\mathscr{F}\times\mathscr{F}\to\mathbb{R}_{0}^{+}$, $\mathcal{S}^{h,\xi}$ is bounded below on $\mathfrak{M}_{1}(\mathscr{F})$ and thus $m:=\inf_{\mathfrak{M}_{1}(\mathscr{F})}\mathcal{S}^{h,\xi}$ exists in $[0,\infty)$, we can choose a minimizing sequence $(\mu_{j})\subset\mathfrak{M}_{1}(\mathscr{F})$ for $\mathcal{S}^{h,\xi}$, so that in particular $m=\lim_{j\to\infty}\mathcal{S}^{h,\xi}(\mu_{j})$. By the duality relation $\operatorname{C}_{0}(\mathscr{F})^{\prime}\cong\mathfrak{M}(\mathscr{F})$ and using that $\mathfrak{M}_{1}(\mathscr{F})$ is convex and closed, the Banach-Alaoglu theorem provides us with a non-relabeled subsequence and a probability measure $\mu\in\mathfrak{M}_{1}(\mathscr{F})$ such that we have $\mu_{j}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\mu$ in $\mathfrak{M}_{1}(\mathscr{F})$. By Lemma 4.1, $\mathcal{S}$ is continuous with respect to weak*-convergence. Now, if (i) $d$ is the Fréchet metric, then $d(\cdot,\rho)=\|\cdot-\rho\|_{\mathfrak{M}(\mathscr{F})}$ is lower semicontinuous with respect to weak*-convergence. On the other hand, if (ii) $d$ is the $p$-Wasserstein metric, then $d$ metrizes weak*-convergence and so, in particular, $d(\cdot,\rho)$ is continuous with respect to weak*-convergence. In both cases, ${\mathcal{S}}^{h,\xi}$ is lower semicontinuous with respect to weak*-convergence. Hence,

### v2

@@ 1. Introduction
The theory of causal fermion systems is a recent approach to fundamental physics (for an introduction to the physical background and applications as well as the mathematical context, we refer the interested reader to the reviews [9, 11], the textbooks [7, 10] or the website [1]). In this approach, spacetime and all structures therein are encoded in a measure $\rho$ on a set of operators on a Hilbert space. The physical equations are formulated via a variational principle for the measure $\rho$, the so-called causal action principle. Causal variational principles evolved as a mathematical generalization of the causal action principle [6, 12] (an introduction to the causal action principle and causal variational principles can be found for example in [10, Chapters 5 and 6]). From the point of view of the calculus of variations, causal variational principles are a class of nonlinear, non-convex variational principles where one minimizes an action $\mathcal{S}$} under variations of a measure $\rho$. One of the objectives of the present paper is to formulate and analyze corresponding flows of measures. Moving from the study of minimizing measures to flows of measures can be understood in analogy to the transition from stationary problems (like for example minimizing the Dirichlet energy) to corresponding evolution equations (like for example the heat flow). In simple terms, our flows can be understood as gradient flows corresponding to causal variational principles. Due to the lack of convexity and smoothness, the formulation of the flow equations as well the proof of existence of solutions are mathematically challenging and seem of general interest in the context of non-smooth and non-convex variational problems.
1.1. Causal variational principles
In order to describe this objective and underlying obstructions in more detail, we begin by recalling the general setting of causal variational principles. For simplicity, we firstly restrict attention to the so-called compact setting; the detailed set-up shall be deferred to Section 2.1 below.
Our starting point is a compact metric space $(\mathscr{F},d)$ and a non-negative function ${{\mathcal{L}}}:\mathscr{F}\times\mathscr{F}\rightarrow\mathbb{R}^{+}_{0}:=[0,\infty)$ (the Lagrangian) which is assumed to be continuous. The corresponding causal action principle then is to
@@ 1.1. Causal variational principles
over the class $\mathfrak{M}_{1}(\mathscr{F})$ of normalized Borel measures on $\mathscr{F}$. Causal variational principles are a class of examples for non-smooth and non-convex variational principles. The existence of solutions of (1.1) is a consequence of the direct method of the Calculus of Variations (see Section 2.1). Most importantly, minimizers $\rho$ satisfy the corresponding Euler-Lagrange equations (EL equations for brevity), and their precise formulation is given in Section 2.1.
Constructing solutions of the EL equations – or physically meaningful approximations thereof – is of central importance in the theory of causal fermion systems in order to get a better understanding of the nature of the physical interactions as described by the causal action principle. Here, abstract existence results are not sufficient, but one needs constructive methods which give insight into the structure of the minimizing measure. By the aforementioned lack of smoothness and non-convexity, this is a non-trivial task in itself. In this regard, a central objective of the general theory is to find a canonical way of how a generic probability measure $\rho_{0}$ can be modified continuously to yield an (approximate) solution of the EL equations. In other words, this corresponds to a meaningful evolution $t\mapsto\varrho(t)$ with $\varrho(0)=\rho_{0}$ such that, for $t\to\infty$, $\varrho(t)$ approaches an (approximate) solution of the EL equations.
1.2. Gradient flows
By the variational nature of the problems considered here, it is natural to consider evolutions driven by the energies or actions given by (1.1). By this we mean that the energies of the solutions are decreasing in time. Heuristically, this can be interpreted as a measure-valued variant of the ordinary differential equation
$\left\{\begin{array}[]{rll}\displaystyle\frac{\operatorname{d}\!}{\operatorname{d}\!t}\varrho(t)\!\!\!&=-\nabla{\mathcal{S}}(\varrho)&\;\text{if~$t>0$}\>,\\[5.0pt] \varrho(0)\!\!\!&=\rho_{0}\>.\end{array}\right$.\ (1.4)
However, for future reference, we remark that (1.4) has to be understood symbolically; in our case and as shall be discussed below, this is due to the lack of smoothness, in turn being a consequence of the non-convexity and non-smoothness of the action $\mathcal{S}$.
By way of comparison, in the more familiar situation of classical Dirichlet energies e.g. on Sobolev spaces, (1.4) reduces to the usual heat equation. The convexity of the underlying energies then allows for useful a priori estimates, finally leading to both existence and regularity assertions for the respective flows. These methods have been refined and extended to many other flow equations, provided that the driving energies are convex.
1.3. Flows for non-convex variational problems
The situation changes drastically if the underlying energies are no longer convex. To the best of our knowledge, there is no unifying theory that yields both existence and decisive statements on the long-time behavior of solutions of the associated gradient flows. To overcome the first issue, we employ a version of De Giorgi’s minimizing movements approach [5, 2, 4] adapted to the present setting; in essence, they can be understood as a method for extending the gradient flow to non-smooth actions on infinite-dimensional spaces.. This construction leads to a flow
@@ 1.3. Flows for non-convex variational problems
with the property that the action given by (1.1) is strictly decreasing along the flow lines. In essence, this is achieved by solving variational problems in discrete time steps which are penalized by the Wasserstein metric, and then pass to a continuous time evolution by use of an Arzelà-Ascoli-type argument. While we describe an analogous penalization procedure by use of the total variation norm, the use of the Wasserstein metric is most suitable here. Indeed, it is the weak*-convergence of probability measures for which compactness can be achieved and the actions (1.1) are lower semicontinuous; the Wasserstein metric, in turn, induces weak*-convergence. We also study the analogous procedure for the total variation norm. In this case, we also get existence of a flow. But the flow has the shortcoming that it potentially gets stuck away from local minima (as will be explained in an example in Section 5). With this in mind, it seems that the Wasserstein distance is the correct metric for the flow of measures we have in mind. We prove that the resulting curves of measures are Hölder continuous (see Section 4.3).
It is an important task to control the long-time behavior of solutions. It is here where the interplay of non-convexity and the weak compactness properties of weak*-convergence necessitate additional arguments. First, it is clear from the arbitrariness of the initial value $\rho_{0}$ that, at best, the curve will converge to an extremal point but not necessarily to a minimizer. fact, by the very definition of the flow, it might get stuck at a critical point of the functional, and by the non-convexity, the latter might be far away from any global minimizer. In the general situation considered here, the situation is even worse: it may happen that the gradient flow does not converge at all. This will be shown in Section 3 in a simple example where the potential is constructed as a downward spiral with increasingly small potential wells (see Figure 1 on page 1). In examples of this type, which may be known to the experts in different scenarios, there is not even a subsequence of times $(t_{k})$ for which the measures converge to a solution of the EL equations.
@@
In order to overcome such difficulties, we also introduce another flow which involves an additional penalization term involving a parameter $\xi>0$. In the case $\xi=0$, we get back the above flow by minimizing movements. In the case $\xi>0$, the additional penalization term gives us a-priori control of the length of the curve (as measured in the Wasserstein distance) in terms of the change of the action (see Section 4.4). This makes it possible to reparametrize the curve, using the action itself as the new parameter. In this way, we can circumvent the difficulty that the flow might get stuck in “plateaus” of the potential for a long time (as shown in Figure 2 on page 2). After the reparametrization, the curve becomes even Lipschitz continuous (see Section 4.4). Moreover, we get control of the long-time behavior of the solutions. Indeed, in the case $\xi>0$ we prove that the resulting curve $\varrho^{\xi}(t)$ does converge (see Section 4.5). The prize to pay is that the limiting measure satisfies the EL equations only approximately. For the error term, we derive a precise a-priori bound which tends to zero as $\xi\searrow 0$. With this in mind, our procedure seems well-suited for the applications in mind. For example, in a numerical study one can choose $\xi$ so small that the error of the approximation is bounded by the numerical errors.
Towards the limiting behavior of the flow as $t\to\infty$, we thus introduce a re-parametrization that accelerates the discrete flow along energy plateaus. This re-parametrization is defined in terms of the underlying action and simultaneously gives bounds on the lengths of the associated curves. As we establish in Section 4.5, the very method only allows for such a construction if…
We also extend our methods and results to the causal action principle for causal variational principles. Our methods and results can be understood more generally from the perspective of non-convex variational problems. Indeed, causal variational principles are model examples of variational principles which, in general, are fully non-convex. The methods to be developed in the present paper provide Hölder continuous flows of measures with these desired properties.
1.4. Structure of the paper
@@ 2.1. Causal variational principles in the compact setting
We let $(\mathscr{F},d)$ be a compact metric space and suppose that the Lagrangian ${\mathcal{L}}\colon\mathscr{F}\times\mathscr{F}\to\mathbb{R}_{0}^{+}$ satisfies the following assumptions:
@@
[…]
@@
${\mathcal{S}}(\rho)=\int_{\mathscr{F}}\operatorname{d}\!\rho(x)\int_{\mathscr{F}}\operatorname{d}\!\rho(y)\>{\mathcal{L}}(x,y)$ (2.1)
@@
$M:=\operatorname{supp}\rho:=\mathscr{F}\setminus\bigcup\big\{U\subset\mathscr{F}\;\text{open}\colon\;\rho(U)=0\big\}\>.$
@@ 2.2. Background facts from optimal transport and metric measure spaces
In our arguments below, we will also make use of the $p$-Wasserstein metric on $\mathscr{F}$ for $1\leqslant p<\infty$. Given a measure $\mathbb{P}\in\mathfrak{M}_{1}(\mathscr{F}\times\mathscr{F})$, for $i\in\{1,2\}$ we denote the projection to the $i^{\text{th}}$ component by $\pi^{i}\colon\mathscr{F}\times\mathscr{F}\ni(x_{1},x_{2})\mapsto x_{i}\in\mathscr{F}$. We let $\pi_{\#}^{i}\mathbb{P}(A):=\mathbb{P}(\pi_{i}^{-1}(A))$ for $A\subset\mathscr{F}$ be the corresponding push-forward of $\mathbb{P}$. As is customary in this context, we then define for $\mu_{1},\mu_{2}\in\mathfrak{M}_{1}(\mathscr{F})$ the class of couplings $\Gamma(\mu_{1},\mu_{2})$ (also referred to as transport plans) by
@@
Here the measures $\pi_{\#}^{i}\mathbb{P}$ are referred to as marginals. Let $1\leqslant p<\infty$. We then define for $\mu,\nu\in\mathfrak{M}_{1}(\mathscr{F})$ the $p$-th Wasserstein metric by
@@
It has the desired marginals $\pi_{\#}^{1}\mathbb{P}=\mu$ and $\pi_{\#}^{2}\mathbb{P}=\nu$. We thus obtain the estimate
$W_{p}(\mu,\nu)^{p}$ $\leqslant\iint_{\mathscr{F}\times\mathscr{F}}d(x,y)^{p}\>d\mathbb{P}(x,y)$
@@
$=\frac{1}{2}\>\text{diam}(\mathscr{F})^{p}\>\|\mu-\nu\|_{\mathfrak{M}(\mathscr{F})}\>.$
@@
In order to prove (2.7), we let $\varepsilon>0$ be arbitrary and choose $\mathbb{P}\in\Gamma(\mu,\nu)$, $\widetilde{\mathbb{P}}\in\Gamma(\nu,\nu)$ such that
@@ 3. An example of a non-smooth, non-convex variational principle
$\dot{\gamma}(t)=-\nabla{\mathcal{S}}|_{\gamma(t)}\qquad\text{and}\qquad\gamma(0)=\Big(r=\frac{1}{5},\varphi=0\Big)\>.$ (3.1)
@@
${\mathcal{S}}(r,\varphi)+\frac{1}{2h}\>d\big((r,\phi),(r^{\prime},\phi^{\prime})\big)^{2}\>,$ (3.2)
where $d$ denotes the Euclidean distance in $\mathbb{R}^{2}$. Indeed, computing the first variation of this action in the Cartesian variables $x=r\cos\phi$ and $y=r\sin\phi$, we obtain the EL equations
$\begin{pmatrix}\partial_{x}{\mathcal{S}}\\ \partial_{y}{\mathcal{S}}\end{pmatrix}+\frac{1}{h}\>\begin{pmatrix}x-x^{\prime}\\ y-y^{\prime}\end{pmatrix}=0\>.$
Assuming that the limit $h\searrow 0$ exists, we obtain the differential equation (3.1). Therefore, the penalized action (3.2) can be regarded as a discrete version of the gradient flow with step size $h$.
@@
Now the corresponding flow equation takes the form
@@ 4.1. The causal action with penalization
$\text{Case~1.}\;\;d=d_{\mathfrak{M}(\mathscr{F})}\qquad\text{or}\qquad\text{Case~2.}\;\;d=W_{p}\>.$ (4.2)
@@
Let $(\mathscr{F},d)$ be a compact metric space and let ${\mathcal{L}}\in\operatorname{C}(\mathscr{F}\times\mathscr{F})$. Then the functional
@@
[…]
@@
Let $\rho,\rho_{1},\rho_{2},...\in\mathfrak{M}_{1}(\mathscr{F})$ be such that $\rho_{j}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\rho$ as $j\to\infty$. Since $\mathcal{F}$ is compact, the Weierstraß  approximation theorem implies that the space
@@
is dense in $\operatorname{C}(\mathscr{F})$. Let $\varepsilon>0$ be arbitrary but fixed. We then find $h\in\operatorname{C}(X\times X)$ of the form $h(x,y)=\sum_{i=1}^{N}h_{i}f_{i}(x)g_{i}(y)$ with $h_{1},...,h_{N}\in\mathbb{R}$ such that $\|{\mathcal{L}}-h\|_{\infty}<\varepsilon$. Therefore,
$\left|\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)-\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)\right|$
$\leqslant\iint_{\mathscr{F}\times\mathscr{F}}|{\mathcal{L}}(x,y)-h(x,y)|\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)$
$\quad\>+\left|\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)-\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)\right|$
$\quad\>+\iint_{\mathscr{F}\times\mathscr{F}}|{\mathcal{L}}(x,y)-h(x,y)|\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)=:\mathrm{I}+\mathrm{II}+\mathrm{III}.$
@@
$\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho_{j}(x)\operatorname{d}\!\rho_{j}(y)$ $=\sum_{i=1}^{N}h_{i}\Big(\int_{\mathscr{F}}f(x)\operatorname{d}\!\rho_{j}(x)\Big)\Big(\int_{\mathscr{F}}g(y)\operatorname{d}\!\rho_{j}(y)\Big)$
@@
$=\iint_{\mathscr{F}\times\mathscr{F}}h(x,y)\operatorname{d}\!\rho(x)\operatorname{d}\!\rho(y)$
@@
Since ${\mathcal{L}}\colon\mathscr{F}\times\mathscr{F}\to\mathbb{R}_{0}^{+}$, $\mathcal{S}^{h,\xi}$ is bounded below on $\mathfrak{M}_{1}(\mathscr{F})$ and thus $m:=\inf_{\mathfrak{M}_{1}(\mathscr{F})}\mathcal{S}^{h,\xi}$ exists in $0,\infty)$, we can choose a minimizing sequence $(\mu_{j})\subset\mathfrak{M}_{1}(\mathscr{F})$ for $\mathcal{S}^{h,\xi}$, so that in particular $m=\lim_{j\to\infty}\mathcal{S}^{h,\xi}(\mu_{j})$. By the duality relation $\operatorname{C}_{0}(\mathscr{F})^{\prime}\cong\mathfrak{M}(\mathscr{F})$ and using that $\mathfrak{M}_{1}(\mathscr{F})$ is convex and closed, the Banach-Alaoglu theorem provides us with a non-relabeled subsequence and a probability measure $\mu\in\mathfrak{M}_{1}(\mathscr{F})$ such that we have $\mu_{j}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\mu$ in $\mathfrak{M}_{1}(\mathscr{F})$. By Lemma 4.1, $\mathcal{S}$ is continuous with respect to weak*-convergence. Now, if (i) $d$ is the Fréchet metric, then $d(\cdot,\rho)=\|\cdot-\rho\|_{\mathfrak{M}(\mathscr{F})}$ is lower semicontinuous with respect to weak*-convergence. On the other hand, if (ii) $d$ is the $p$-Wasserstein metric, then $d$ metrizes weak*-convergence and so, in particular, $d(\cdot,\rho)$ is continuous with respect to weak*-convergence. In both cases, ${\mathcal{S}}^{h,\xi}$ is lower semicontinuous with respect to weak*-convergence. Hence,